AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 19 of 33  ELOG logo
ID Date Author Subject Text Attachments
  304   Sun Feb 27 12:49:44 2022 TD RSSSunday 27 February 07:00-15:00
13.50 Cryostat set point 120K
 6x 
  303   Sun Feb 27 12:00:13 2022 RSS TDCurrent statusRing current ~7.7e7 particles/s

Merger, Tape server (WS3):
1414 kB/s writing ok
 13-xy-2022-02-27_13-01-36.png13-yrate-2022-02-27_13-01-06.png13-xrate-Screenshot_from_2022-02-27_13-00-50.png 
  302   Sun Feb 27 11:00:08 2022 RSS TDCurrent status12:00: 95K-->120K

12:07: 120K --> 95K

Ring current ~6e7 particles/s
 12-xy-2022-02-27_12-16-19.png12-yrate-2022-02-27_12-15-42.png12-xrate-2022-02-27_12-15-03.png 
  301   Sun Feb 27 10:00:07 2022 RSS TDCurrent statusRing current ~9e7 particles/s

Merger, Tape server (WS3):
1414 kB/s writing ok
 11-xy-2022-02-27_11-01-22.png11-ey-rate2022-02-27_11-00-46.png11-ex-rate-2022-02-27_11-00-27.png 
  300   Sun Feb 27 09:00:09 2022 RSS TDCurrent status10:00: 95K-->120K

10:08: 120K --> 95K

Ring current ~7e7 particles/s
 7x 
  299   Sun Feb 27 08:00:23 2022 RSS TDCurrent statusRing current ~7.5e7 particles/s

Merger, Tape server (WS3):
1414 kB/s writing ok
 9-xy-2022-02-27_09-00-57.png9-ey-2022-02-27_09-00-03.png9-ex-2022-02-27_08-59-48.png 
  298   Sun Feb 27 06:58:51 2022 RSS TDCurrent status7:59: 95K-->120K

8:06: 120K --> 95K

Ring current ~7e7 particles/s
 8-xy-2022-02-27_08-10-15.png8-ey-2022-02-27_08-09-46.png8-ex-2022-02-27_08-09-29.pngScreenshot_from_2022-02-27_07-16-30.png 
  297   Sun Feb 27 06:23:42 2022 RSS TDCurrent statusRing current ~6e7 particles/s

Merger, Tape server (WS3):
1414 kB/s writing ok
 6x 
  296   Sun Feb 27 05:45:53 2022 David, Pierre-MichellCurrent status6:05: 95K-->120K

6:12: 120K --> 95K

Ring current 8e7 particles/s
 2022-02-27_06-50-09exrat.png2022-02-27_06-48-23eytot.png2022-02-27_06-48-05extot.png2022-02-27_06-47-47xytot.png 
  295   Sun Feb 27 04:54:46 2022 David, Pierre-MichellCurrent statusRing current 8e7 particles/s

Merger, Tape server (WS3):
1664 kB/s writing ok
 2022-02-27_05-56-59exra.png2022-02-27_05-56-31exto.png2022-02-27_05-56-13xytot.png 
  294   Sun Feb 27 03:49:36 2022 David, Pierre-MichellCurrent status04:06 85K -->120K : 04:13 120K -->95K

current ring 8e7 p/s

Merger, Tape server (WS3):
 2022-02-27_04-53-12xytot.png2022-02-27_04-52-51exrst.png2022-02-27_04-52-31extot.png 
  293   Sun Feb 27 02:38:52 2022 David, Pierre-MichellCurrent statusWe have decreased the temperatur> 95K
--> 90K. At 95K the target was stable,
but let's try to increase the density
 2022-02-27_03-43-36aida.png2022-02-27_03-42-55xyto.png2022-02-27_03-42-14exra.png2022-02-27_03-41-49extot.png 
  292   Sun Feb 27 01:46:56 2022 David, Pierre-MichellCurrent statusSo far looks like with a temperature of
95K the target is maintained with a reasonable
density throughout long periods of time.
 2022-02-27_02-50-15exrate.png2022-02-27_02-49-47extot.png2022-02-27_02-49-28xytot.png 
  291   Sun Feb 27 00:53:24 2022 David, Pierre-MichellCurrent statusWe have recovered the beam at 01:49, after
such a long time we had no target so we increased
from 80K to 120K at 01:52. 02:01 update 120K-->95K
 7x 
  290   Sun Feb 27 00:07:15 2022 David, Pierre-MichellNo beam, valve closedAt 00:37 we lost the beam due to a high
pressure burst in the tarjet which resulted
in a valve closed at YR09VV...
  
  289   Sat Feb 26 23:20:13 2022 David, Pierre-MichellUpdate statusIn this last hour we have been playing
with the cryostat temperature as dictated
by Nikos, in order to have the target working
 7x 
  288   Sat Feb 26 22:22:27 2022 David, Pierre-MichellStarting to take dataAt 23:22 we started recording data under
the new conditions.

Merger, Tape server (WS3):
 11x 
  287   Sat Feb 26 04:36:38 2022 RSS TDSaturday 26 February 07:00-15:0005:36 E4 pressure c. 1.7e-9mbar *not* 3e-10mbar
as reported earlier
      See attachments 1 & 2
 27x 
  286   Fri Feb 25 19:20:22 2022 MS, PJWcheck'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
rabbit decoder status'  no changes
 11x 
  285   Fri Feb 25 17:48:22 2022 MS, PJWcheck'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
rabbit decoder status' - no changes
 11x 
ELOG V3.1.4-unknown